# MERba: Multi-Receptive Field MambaVision for Micro-Expression Recognition

### v1

@@ Abstract
Micro-expressions (MEs) are brief, involuntary facial movements that reveal genuine emotions, holding significant potential in psychological diagnosis and criminal investigations. Despite notable advances in automatic ME recognition (MER), existing methods still struggle to jointly capture localized muscle activations and global facial dependencies, both critical for recognizing subtle emotional cues. To tackle this challenge, we propose MERba, a novel multi-receptive field architecture tailored for MER. MERba introduces a series of Local-Global Feature Integration stages, where fine-grained motion features are first extracted by local extractors containing MambaVision Mixers within non-overlapping windows, and then global dependencies across these regions are modeled via lightweight self-attention layers. This hierarchical design enables a progressive transition from localized perception to holistic facial understanding. Furthermore, we introduce an asymmetric multi-scanning strategy to eliminate redundant scanning directions and enhance local spatial perception. To address the high inter-class similarity among negative MEs, we introduce a Dual-Granularity Classification Module that decouples the recognition process into a coarse-to-fine paradigm. Experiments on two benchmark MER datasets demonstrate that MERba outperforms existing methods, with ablation studies confirming the effectiveness of each proposed component.
@@ I Introduction
Recently, the Mamba architecture [11], based on State Space Models (SSMs) [12], has demonstrated outstanding performance across various computer vision tasks [13, 14], particularly favored for its efficient linear time complexity, strong contextual awareness, and simplified network structure. Building on this foundation, MambaVision [15] extends Mamba by integrating Transformer blocks, forming a hybrid structure that aims to balance computational efficiency with long-range spatial dependency modeling. However, directly applying MambaVision to MER tasks is suboptimal, which fails to focus sufficiently on localized facial regions where MEs typically manifest as isolated Action Units (AUs) [16] and neglects fine-grained inter-region dependencies. Indeed, our empirical results show that the baseline MambaVision model underperforms significantly on MER benchmarks, indicating its limited capacity to capture subtle, region-specific motion dynamics.
To tackle this challenge, we propose MERba, a multi-receptive field architecture specifically designed for MER, incorporating the strengths of Mamba and Transformer in a task-adaptive manner. Specifically, MERba integrates fine-grained feature extraction with global dependency modeling through multiple ME Local-Global Feature Integration (LGFI) Stages. In each LGFI stage, instead of scanning the entire image, we define multiple non-overlapping local windows, within which MERba local extractors containing MambaVision Mixers (Mixers) operate independently to extract detailed ME motion cues. Considering emotional muscle movements span multiple facial areas, we employ global self-attention blocks after local extractors, where even a few Transformer layers per stage are sufficient to effectively model dependencies across local windows. Moreover, stage-wise downsampling expands the receptive field progressively, enabling a hierarchical transition from localized perception to full-face understanding.
@@
Additionally, to address the fine-grained nature of emotion recognition, especially in distinguishing subtle negative emotions like anger and disgust, we introduce a Dual-Granularity Classification Module (DGCM). This module uses two parallel classification heads: a coarse classifier that determines emotional polarity (e.g., positive or negative) and a fine classifier that resolves more nuanced emotional categories. This coarse-to-fine paradigm improves the model’s discriminability and robustness, especially under the imbalanced conditions common in ME datasets.
@@
  • We propose MERba, a multi-receptive field architecture for MER, which integrates local extractors based on MambaVision Mixers with global self-attention mechanisms, facilitating fine-grained movement learning and global dependency modeling.
  • We design an asymmetric multi-scanning strategy to eliminate redundant scanning directions and enhance local spatial perception.
  • We introduce a selectively activated Dual-Granularity Classification Module that enhances fine-grained emotion recognition via a coarse-to-fine classification paradigm.
  • Our approach achieves state-of-the-art performance on multiple benchmark ME datasets, with ablation studies validating the effectiveness of each proposed component.
@@ II-A Automatic Micro-Expression Recognition
Early automatic MER methods primarily relied on hand-crafted features, such as LBP-TOP [19] and Bi-WOOF [20]. Although these methods achieved certain success, their performance was limited by the complexity of feature design and adaptability to data variations. With the rise of deep learning, many researchers turned to CNNs for ME feature extraction. Notably, Liong et al. [5] employed a shallow three-stream CNN structure to separately process features from optical flow and optical strain in three dimensions. MERSiamC3D [6] and ME-PLAN [21] both utilized 3D-CNNs for spatiotemporal feature extraction from ME frame sequences. Recently, ViTs have gained traction in MER. Zhang et al. [22] completely abandoned traditional CNNs, relying solely on ViTs and LSTMs to achieve superior performance. Wang et al. [10] proposed HTNet, which stacks multiple hierarchical Transformer layers for feature extraction at different scales. In contrast to these approaches, our method utilizes SSMs for local feature extraction, combined with self-attention mechanisms to integrate global dependencies, enabling comprehensive modeling of fine-grained ME features across multiple receptive fields.
II-B Scanning in Mamba
To enhance the spatial perception of Mamba, many researchers have focused on the design of scanning methods. ViM [13] treated the image as 2D patch sequence and applied bi-directional scanning to capture global contextual information. VMamba [14] scanned bi-directionally along both horizontal and vertical axes, further enriching spatial information extraction. LocalMamba [17] not only performed bi-directional scanning globally but also within small local windows, and customized the scanning direction for each layer. PlainMamba [18] used continuous 2D scanning, maintaining the spatial continuity of adjacent tokens when transitioning between rows or columns. In contrast to above mentioned representative scanning methods, the asymmetric scanning strategy proposed in this paper fully considers the inherent properties of the human face and effectively models spatial relationships for MER tasks while avoiding redundancy.
@@ II-B Scanning in Mamba
Stage 1 Conv $N_{1}=3$ $56\times 56\times 128$ $28\times 28\times 256$
@@
Stage 4 MERba Local Extractor $N_{4}=4$ $7\times 7$ $7\times 7\times 1024$ $1\times 1\times 1024$
@@ III-C Dual-Granularity Classification Module
We introduce DGCM to handle high inter-class similarity in MER. Specifically, unlike traditional approaches that rely on a single linear classification head with output dimension equal to the total number of emotion classes, DGCM adopts a coarse-to-fine classification paradigm via two parallel linear classifiers: a coarse-grained head $h_{\text{coarse}}$ and a fine-grained head $h_{\text{fine}}$. $h_{\text{coarse}}$ predicts the emotional polarity of the input ME sample among four broad categories: negative, positive, surprise, and others (contempt). When a sample is predicted as negative, $h_{\text{fine}}$ is activated to further disambiguate among four easily-confused negative emotions: anger, disgust, fear, and sadness.
During training, the outputs $\hat{y}_{\text{coarse}}$ and $\hat{y}_{\text{fine}}$ from both classification heads are used to compute two cross-entropy losses: $\mathcal{L}_{\text{coarse}}$ and $\mathcal{L}_{\text{fine}}$, respectively. The total training loss $\mathcal{L}_{\text{cls}}$ is progressively weighted to balance learning dynamics:
@@
TABLE II: Sample distribution of DFME-public dataset.1
@@ IV Experiments
In this section, we first provide an overview of the datasets and evaluation metrics used in our experiments. We then detail the implementation settings and present the results of our method. Finally, we conduct an ablation study to analyze the contributions of each component in our model.
@@ IV-A Datasets and Evaluation Metrics
To ensure a comprehensive and fair comparison, we conducted experiments on two representative MER datasets, DFME-public (7 classes) and 3DB-Combined (3 classes). DFME-public refers to the publicly available portion of the DFME dataset [24], subject-independently divided into a fixed training set and two test sets. The model was trained using the entire training set and inference was conducted separately on Test A and Test B, using Unweighted F1-score (UF1), Unweighted Average Recall (UAR) and Accuracy (ACC) as evaluation metrics following [25]. For 3DB-Combined, which is a composite dataset of SMIC-HS [26], CASME II [27], and SAMM [28], we employed Leave-One-Subject-Out (LOSO) cross-validation with UF1 and UAR as evaluation metrics in line with MEGC2019 protocols [29]. Detailed sample distributions of both datasets are provided in Table II and III.
@@ IV-B Implementation Settings
For all datasets, we first apply face detection, alignment, and cropping, following the same procedure as [24]. Subsequently, all images are resized to a uniform resolution of $224\times 224$. For the SMIC-HS dataset, which lacks apex frame labels, we employ the apex frame detection method from [21] to generate the labels. During training, the only data augmentation applied is random horizontal flipping. The optimizer used is AdamW, with the learning rate warmed up over the first 5 epochs to $45e-$ 4, followed by a cosine decay scheduler, with a cooldown period over the last 10 epochs. The weight decay is set to 0.05, and the classification loss function is cross-entropy loss. The total number of training epochs is 200 for the DFME-public dataset. For the 3DB-combined dataset, the number of epochs per subject is 70. Early stopping is employed during training to prevent overfitting, and the dropout rate is set to 0.1. All experiments are conducted using the PyTorch framework, with training and inference performed on a single NVIDIA A100 Tensor Core GPU, and the training batch size is set to 128.
TABLE IV: Comparison of experimental results with other methods on DFME-public dataset.
@@
IV-C Experimental Results
TABLE V: Experimental results compared with state-of-the-art methods on 3DB-Combined.
  MER Methods Backbone Combined SMIC-HS CASME II SAMM
UF1 UAR UF1 UAR UF1 UAR UF1 UAR
@@ IV-C Experimental Results
    MambaVision-B is reproduced with the same parameters as MERba.
IV-C1 DFME-public dataset
DFME-public test sets take FeatRef [7] as the baseline, and the results of top two teams are publicly available based on the official leaderboard [25]. We have also reproduced HTNet [10] and MambaVision-B [15] for a more comprehensive comparison. As shown in Table IV, on both Test Set A and B, MERba with DGCM outperforms all state-of-the-art (SOTA) methods, demonstrating superior performance in the challenging fine-grained 7-class MER task.
IV-C2 3DB-Combined datasets
The results are compared across a range of baseline methods, spanning hand-crafted [19, 20], 2D-CNN [5, 7], 3D-CNN [21], and Transformer-based [22, 9, 10] approaches. MambaVision-B [15] and our MERba represent the latest models using SSM+ViT backbones. As shown in Table V, our method achieves SOTA results on SMIC-HS, SAMM, and the 3DB-combined dataset. On CASME II, it performs on par with the best-performing method (HTNet), trailing by no more than 0.5% in both UF1 and UAR. Notably, on the SAMM dataset, which has a more diverse ethnic groups, our UF1 is 5.7% higher than that of HTNet. Overall, these results highlight the robustness and competitiveness of our approach.
TABLE VI: Ablation study for Local-Global Feature Integration on DFME-public Test B.
@@ IV-C1 DFME-public dataset
TABLE VII: Ablation study for Asymmetric Multi-Scanning on DFME-public Test B.
@@
    $a,b,c,d$ are consistent with the corresponding direction in Figure 2. The subscripts ”$bi$” or ”$sy$” respectively represent the scanning directions exhibiting a bi-directional or horizontal symmetric relationship with the original direction $a$ or $b$.
@@ IV-D1 Local-Global Feature Integration
We perform an ablation study to assess the effect of our proposed LGFI stages on the DFME-public Test B dataset, as shown in Table VI. On the one hand, when self-attention is applied only within the local window like the extractor, the model performs poorly, struggling to capture global dependencies between different facial regions during ME occurrences. On the other hand, when we discard the local window (i.e., both extractors and self-attention are applied to the entire feature map), the model achieves a UF1 of 0.3840 and a UAR of 0.3814. However, it still lacks the fine-grained spatial learning necessary for detecting subtle facial movements. The best performance is achieved by combining local extractors with global self-attention, demonstrating that associating local feature extraction with global dependency modeling provides a substantial improvement in MER performance.

### v2

@@ Abstract
Micro-expressions (MEs) are brief, involuntary facial movements that reveal genuine emotions, offering valuable insights for psychological assessment and criminal investigations. Despite significant progress in automatic ME recognition (MER), existing methods still struggle to simultaneously capture localized muscle activations and global facial dependencies, both essential for decoding subtle emotional cues. To address this challenge, we propose MERba, a hierarchical multi-receptive field architecture specially designed for MER, which incorporates a series of Local-Global Feature Integration stages. Within each stage, detailed intra-window motion patterns are captured using MERba Local Extractors, which integrate MambaVision Mixers with a tailored asymmetric multi-scanning strategy to enhance local spatial sensitivity. These localized features are then aggregated through lightweight self-attention layers that explicitly model inter-window relationships, enabling effective global context construction. Furthermore, to mitigate the challenge of high inter-class similarity among negative MEs, we introduce a Dual-Granularity Classification Module that decomposes the recognition task into a coarse-to-fine paradigm. Extensive experiments on three benchmark datasets demonstrate that MERba consistently outperforms existing methods, with ablation studies confirming the effectiveness of each proposed component.
@@ I Introduction
Recently, the Mamba architecture [11], based on State Space Models (SSMs) [12], has demonstrated outstanding performance across various computer vision tasks [13, 14], particularly favored for its efficient linear time complexity, strong contextual awareness, and simplified network structure. Building on this foundation, MambaVision [15] extends Mamba by integrating Transformer blocks, forming a hybrid structure that aims to balance computational efficiency with long-range spatial dependency modeling. However, directly applying MambaVision to MER tasks is suboptimal, which fails to focus sufficiently on localized facial regions where MEs typically manifest as isolated Action Units (AUs) [16] and neglects fine-grained inter-region dependencies. Indeed, our empirical results (as in Table IV-VI) show that the baseline MambaVision model underperforms significantly on MER benchmarks, indicating its limited capacity to capture subtle, region-specific motion dynamics.
To tackle the above challenges, we propose MERba, a hierarchical multi-receptive field architecture specifically designed for MER, incorporating the strengths of Mamba and Transformer in a task-adaptive manner. Specifically, MERba integrates fine-grained local feature extraction with global dependency modeling through multiple ME Local-Global Feature Integration (LGFI) Stages. In each LGFI stage, instead of scanning the entire image, we define segmented non-overlapping local windows, within which MERba local extractors equipped with MambaVision Mixers (Mixers) operate independently to extract intra-window ME motion details. However, local motion perception remains fragmented and lacks contextual relevance. To address this, we employ global self-attention blocks to aggregate these localized features, explicitly modeling inter-window relationships to establish a holistic global context, thereby obtaining meaningful representations of global emotional movements. Furthermore, we employ progressive downsampling between stages to achieve systematic expansion of the local window receptive field, allowing a hierarchical transition from localized motion perception to comprehensive facial expression understanding.
@@
Additionally, to address the challenge of high semantic similarity and inter-class confusion among negative emotions (e.g., anger and disgust), we introduce a Dual-Granularity Classification Module (DGCM). This module uses two parallel classification heads: a coarse classifier that determines emotional polarity (e.g., positive or negative) and a fine classifier that resolves more nuanced emotional categories. This coarse-to-fine paradigm improves the model’s discriminability and robustness, especially under the imbalanced conditions common in ME datasets.
@@
  • We propose MERba, a novel multi-receptive field architecture for MER, which introduces hierarchical Local-Global Feature Integration stages to enable a progressive transition from detailed localized motion perception to holistic facial understanding.
  • We design an effective local-to-global feature modeling mechanism, where MambaVision Mixers combined with a specialized asymmetric multi-scanning strategy are used to extract fine-grained motion features within non-overlapping windows, and lightweight self-attention layers are employed to capture global dependencies across localized facial regions.
  • We introduce a selectively activated Dual-Granularity Classification Module that decomposes the recognition task into a coarse-to-fine paradigm, effectively reducing inter-class confusion among negative MEs.
  • Our approach achieves state-of-the-art (SOTA) performance on multiple benchmark ME datasets, with ablation studies validating the effectiveness of each proposed component.
@@ II-A Automatic Micro-Expression Recognition
Early automatic MER methods primarily relied on hand-crafted features, such as LBP-TOP [19], Bi-WOOF [20], and MDMO [21]. Although these methods achieved certain success, their performance was limited by the complexity of feature design and adaptability to data variations. With the rise of deep learning, many researchers turned to CNNs for ME feature extraction. Liong et al. [5] employed a shallow three-stream CNN structure to separately process features from optical flow and optical strain in three dimensions. Zhou et al. [7] used a two-stream Inception network to learn expression-shared features, followed by Softmax attention for learning expression-specific features and feature fusion. Wang et al. [22] proposed TLCNN, which combines Deep CNN for spatial features and LSTM for temporal modeling. MERSiamC3D [6] and ME-PLAN [23] both utilized 3D-CNNs for spatiotemporal feature extraction from ME frame sequences. Recently, ViTs have gained traction in MER. Zhang et al. [24] completely abandoned traditional CNNs, relying solely on ViTs and LSTMs to achieve superior performance. Wang et al. [10] proposed HTNet, which stacks multiple hierarchical Transformer layers for feature extraction at different scales. In contrast to these approaches, our method utilizes SSMs for local feature extraction, combined with self-attention mechanisms to integrate global dependencies, enabling comprehensive modeling of fine-grained ME features across multiple receptive fields.
@@
Stage 1 Conv $N_{1}=3$ - $56\times 56\times 128$ $28\times 28\times 256$
@@
Stage 4 MERba Local Extractor $N_{4}=4$ $\times 7 7×7$$\times 7\times 1024$ $1\times 1\times 1024$
@@
II-B Scanning in Mamba
To enhance the spatial perception of Mamba, many researchers have focused on the design of scanning methods. ViM [13] treated the image as 2D patch sequence and applied bi-directional scanning to capture global contextual information. VMamba [14] scanned bi-directionally along both horizontal and vertical axes, further enriching spatial information extraction. LocalMamba [17] not only performed bi-directional scanning globally but also within small local windows, and customized the scanning direction for each layer. PlainMamba [18] used continuous 2D scanning, maintaining the spatial continuity of adjacent tokens when transitioning between rows or columns. In contrast to above mentioned representative scanning methods, the asymmetric multi-scanning strategy proposed in this paper fully considers the inherent properties of the human face and effectively models spatial relationships for MER tasks while avoiding redundancy.
@@ III-C Dual-Granularity Classification Module
We introduce a Dual-Granularity Classification Module (DGCM) to handle high inter-class similarity in fine-grained MER scenarios. Specifically, unlike conventional approaches that rely on a single linear classification head with an output dimension equal to the total number of emotion classes, DGCM adopts a coarse-to-fine classification paradigm via two parallel linear classifiers: a coarse-grained head $h_{\text{coarse}}$ and a fine-grained head $h_{\text{fine}}$. $h_{\text{coarse}}$ first categorizes each ME sample into one of several broad affective groups (e.g., negative, positive, surprise, and optionally others or contempt). When a sample is predicted as negative, $h_{\text{fine}}$ is activated to further disambiguate among multiple visually similar negative emotions, such as anger, disgust, fear, and sadness. This coarse-to-fine mechanism encourages the model to focus its representational capacity where fine distinctions are most needed, while maintaining robustness and generalizability.
To effectively optimize this hierarchical prediction process, we introduce a two-branch training objective with progressive weighting between the coarse and fine levels. Specifically, during training, the outputs $\hat{y}_{\text{coarse}}$ and $\hat{y}_{\text{fine}}$ from both classification heads are used to compute two cross-entropy losses: $\mathcal{L}_{\text{coarse}}$ and $\mathcal{L}_{\text{fine}}$, respectively. The total training loss $\mathcal{L}_{\text{cls}}$ is progressively weighted to balance learning dynamics:
@@
TABLE II: Sample distribution of DFME-public and MMEW datasets.1
@@
TABLE IV: Experimental results compared with SOTA methods on 3DB-Combined datasets.
  MER Methods Backbone Combined SMIC-HS CASME II SAMM
UF1 UAR UF1 UAR UF1 UAR UF1 UAR
  • 1
    MambaVision-B is reproduced with the same parameters as MERba.
@@ IV Experiments
In this section, we first provide an overview of the datasets and evaluation metrics used in our experiments. We then detail the implementation settings and present the results of our method. Finally, we conduct an ablation study to analyze the contributions of each component in our model, and include a case study to provide visual explanations for how the model captures key emotional cues.
@@ IV-A Datasets and Evaluation Metrics
To ensure a comprehensive and fair comparison, we conducted experiments on three representative MER datasets, 3DB-Combined (3 classes), DFME-public (7 classes), and MMEW (6 classes). Detailed sample distributions of the three datasets involved in our experiments are provided in Table II and III.
  • 3DB-Combined is a composite dataset of SMIC-HS [26], CASME II [27], and SAMM [28]. We employed Leave-One-Subject-Out (LOSO) cross-validation with Unweighted F1-score (UF1) and Unweighted Average Recall (UAR) as evaluation metrics in line with MEGC2019 protocols [29].
  • DFME-public refers to the publicly available portion of the DFME dataset [30], subject-independently divided into a fixed training set and two test sets. The model was trained using the entire training set and inference was conducted separately on Test A and Test B. Following [31], we used UF1, UAR, and Accuracy (ACC) as evaluation metrics.
  • MMEW [1] contains 300 ME samples across 7 emotion categories. To align with the experimental protocol in the original paper, we excluded the others category. The remaining 234 samples, covering 6 emotion classes, were used for subject-independent 5-fold cross-validation. ACC was adopted as the evaluation metric, consistent with the benchmark setting.
@@ IV-B Implementation Settings
For all datasets, we first apply face detection, alignment, and cropping, following the same procedure as [30]. Subsequently, all images are resized to a uniform resolution of $224\times 224$. For the SMIC-HS dataset, which lacks apex frame labels, we employ the apex frame detection method from [23] to generate the labels. During training, the only data augmentation applied is random horizontal flipping. The optimizer used is AdamW, with the learning rate warmed up over the first 5 epochs to 5e45e-4, followed by a cosine decay scheduler, with a cooldown period over the last 10 epochs. The weight decay is set to 0.05, and the classification loss function is cross-entropy loss. The total training epochs are 200 for DFME-public and each fold of MMEW. For 3DB-combined, the number of epochs per subject is 70. Early stopping is employed during training to prevent overfitting, and the dropout rate is set to 0.1. All experiments are conducted using the PyTorch framework, with training and inference performed on a single NVIDIA A100 Tensor Core GPU, and the training batch size is set to 128.
IV-C Experimental Results
IV-C1 3DB-Combined
The results are compared across a range of baseline methods, spanning hand-crafted [19, 20], 2D-CNN [5, 7], 3D-CNN [23], and Transformer-based [24, 9, 10] approaches. MambaVision-B [15] and our MERba represent the latest models using SSM+ViT backbones. Since the 3DB-Combined dataset adopts a 3-class protocol without fine-grained negative emotion labels, we report results using MERba without the DGCM module. As shown in Table IV, our method achieves SOTA results on SMIC-HS, SAMM, and the overall 3DB-combined dataset. On CASME II, MERba performs competitively with HTNet, trailing by less than 0.5% in both UF1 and UAR. Notably, on the SAMM dataset, which has a more diverse ethnic groups, our model surpasses HTNet by a large margin of 5.7% UF1, demonstrating stronger generalization across demographic variation. Moreover, MERba achieves this performance with only 101.21M parameters—over four times smaller than HTNet’s 438.51M—highlighting its superior parameter efficiency.
TABLE V: Comparison with SOTA methods on DFME-public.
@@ IV-C Experimental Results
IV-C2 DFME-public
DFME-public test sets take FeatRef [7] as the baseline, and we further include the results of top two teams from the DFME Challenge at CCAC 2024, which are publicly available on the official leaderboard [31]. We have also reproduced HTNet [10] and MambaVision-B [15] for a more comprehensive comparison. As shown in Table V, on both Test Set A and B, MERba with DGCM outperforms all SOTA methods, demonstrating superior performance in the challenging fine-grained 7-class MER task.
TABLE VI: Comparison with SOTA methods on MMEW.
  MER Methods Backbone ACC(%)
LBP-TOP (2014)[19] Hand-crafted 38.9
MDMO (2015)[21] Hand-crafted 65.9
TLCNN (2018)[22] CNN+LSTM 69.4
LD-FMERN (2023)[32] 2D-CNN 71.7
HTNet1  (2024)[10] ViT 71.8
MA2MI (2024)[33] transfer learning 75.2
MambaVision-B1  (2024) [15] SSM+ViT 70.1
MERba-base2  (ours) SSM+ViT 72.6
MERba-DGCM (ours) SSM+ViT 75.2
@@ IV-C2 DFME-public
  • 2
    MERba-base denotes replacing DGCM with a single linear head which predicts all six emotion categories.
IV-C3 MMEW
As shown in Table VI, MERba outperforms all baseline models based on hand-crafted features, CNNs, and Transformers on the MMEW dataset. MERba with DGCM achieves the highest accuracy of 75.2%, performing on par with MA2MI [33], which denotes a SOTA model employing transfer learning. Notably, MA2MI leverages extensive macro-expression data for pretraining, while our MERba is trained solely on ME samples, yet reaches top-level performance. This highlights how our architecture effectively adapts to the MER task and captures fine-grained emotional distinctions, even with limited supervision.
TABLE VII: Ablation study for Local-Global Feature Integration on DFME-public Test B.
@@ IV-C3 MMEW
TABLE VIII: Ablation study for Asymmetric Multi-Scanning on DFME-public Test B.
@@
    $a,b,c,d$ are consistent with the corresponding direction in Figure 2. The subscripts ”bibi” or ”sysy” respectively represent the scanning directions exhibiting a bi-directional or horizontal symmetric relationship with the original direction $a$ or $b$.
@@ IV-D1 Local-Global Feature Integration
We perform an ablation study to assess the effect of our proposed LGFI stages on the DFME-public Test B dataset, as shown in Table VII. On the one hand, when self-attention is applied only within the local window like the extractors, the model performs poorly, struggling to capture global dependencies between different facial regions during ME occurrences. On the other hand, when we discard the local window (i.e., both extractors and self-attention are applied to the entire feature map), the model achieves a UF1 of 0.3840 and a UAR of 0.3814. However, it still lacks the fine-grained spatial learning necessary for detecting subtle facial movements. The best performance is achieved by combining local extractors with global self-attention, demonstrating that associating local feature extraction with global dependency modeling provides a substantial improvement in MER performance.